OpenStudy (angiepangie0726):

Help please?? Fan and medal ~ Solve 2x^2 + 16x + 27 = 0 by completing the square.

OpenStudy (sshayer):

\[x^2+8x=-\frac{ 27 }{ 2 }\] adding both sides \[\left( \frac{ 8 }{ 2 } \right)^2or~16\] \[x^2+8x+16=16-\frac{ 27 }{ 2 }=\frac{ 5 }{ 2 }\] \[\left( x+4 \right)^2=\frac{ 5 }{ 2 }\] \[x+4=\pm \frac{ 5 }{ 2 }\] \[x=-4 \pm \frac{ 5 }{ 2 }\] find both values of x

OpenStudy (angiepangie0726):

Thank you so much!! I have one more question, mind if I open another on and ask you??

OpenStudy (angiepangie0726):

one*